OpenStudy (anonymous):

The interest on $6,000 at 6 percent compounded semiannually for 8 years is

OpenStudy (anonymous):

\[6000(1.06)^8\] and a calculator

OpenStudy (campbell_st):

is the interest rate per year or per half year...?

OpenStudy (anonymous):

oops semi anually good catch try \[6000(1.03)^{16}\]

OpenStudy (anonymous):

how did you come up with 1.06?

OpenStudy (campbell_st):

have a look at the 2nd solution posted by satellite...

OpenStudy (anonymous):

im confused

OpenStudy (campbell_st):

ok... the formula for compound interest is A = P(1 + r/100)^n A = future value, P = Principal, r is the interest rate n is the number of time periods.. you question says the interest is compounded half yearly so n = 2 x 8 or 16 there are 16 half years in 8 full years r = 6 or 0.06 if the interest rate is 6% per annum then it becomes 3% per half year. so in your question you have A = 6000(1 +3/100)^16 does this make sense?

OpenStudy (anonymous):

looking on the table I get a 1.6047. If I multiply that by 6000 I get a total of $9,628.20.

OpenStudy (campbell_st):

and that answer can be obtained by using the formula above...

OpenStudy (anonymous):

how?

OpenStudy (campbell_st):

do you have a calculator...?

OpenStudy (anonymous):

no

OpenStudy (campbell_st):

here is a link to an online calculator... I'll assume your computer doesn't have one. just enter, as its written below \[6000 \times ( 1 + 3 \div 100)^{16} = \]

OpenStudy (anonymous):

yeah I have one on my computer, thought you meant a hand calculator..but what the link it should come in handy

OpenStudy (anonymous):

\(6000(1.03)^{16}\) try this http://www.wolframalpha.com/input/?i=6000%281.03%29^16